State with reasons whether the following statements are True or False.
i) The original signal can never be reproduced at the receiver end when frequency
modulation technique is used.
1
Expert's answer
2019-04-19T10:32:57-0400

Frequency modulation was created to vary the carrier frequency using the information signal and it allows reproducing original information signals:

- Original signal: "V_m(t),"

- Carrier signal: "V_C(t)=V_{C\text{max}}\text{sin}(2\pi f_C+\phi),"

- Frequency modulated signal: "V_{FM}=V_{C\text{max}}\text{sin}(2\pi [f_C+\frac{\Delta f}{V_{m.\text{max}}}V_m(t)]+\phi)."

So in the FM signal we see the original signal. The modulated signal then is transmitted to the receiver. In the receiver this signal passes through the FM demodulator, low pass filter and finally the original signal is reproduced.